OpenStudy (anonymous):

QUADRATIC EQUATIONS Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. 0 = x2 + 5x + 6 0 = x2 + 4x + 4 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 4x + 4 different from y = x2 + 5x + 6?

OpenStudy (cwrw238):

0 = x2 + 5x + 6 (x + 2)(x + 3) = 0 x = -2 or -3 0 = x2 + 4x + 4 (x + 2^2 = 0 x = -2

OpenStudy (cwrw238):

first equation will open upwards and will intersect the x-axis at x = -2 and x=-3 second equation has a double root = -2 it will open upwards and the bottom of graph will touch x-axis at x = -2

OpenStudy (anonymous):

That is part two right?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

Thank you!!

OpenStudy (cwrw238):

yw have u got welsh ancestry megan? ur first and last names are welsh

OpenStudy (cwrw238):

or are u welsh? - hope u don't mind me asking

OpenStudy (anonymous):

Yes i am welsh

OpenStudy (cwrw238):

good - there at least 2 welsh girls on the sire today